Solve the equation. Round your answer to three decimal places. 55) logx (89.5) = 3.6 56) 5X = 23 57) 2(x - 2) = 18Use Cramers rule to solve the system. 62) 4x + 3y = 26 4x + 2y = 20 Find the determinant of the given matrix, [-1-3-4 63) 5
Answer #1

& log, 89.5 = 3.6 23:6 - 89.5 a= ( 89.5836 [x = 3.48476 56 59 =23 log 23 - a 1851.908197 2C2-2) 18 2942-18 on - 18 a = 72 2 =Cramer rak WW+39=26 Ua+2y=20 126 B- A 4 2 1 IAI= 8-12=-4 x 13 52-608-2 y. 32. Bo-lov. 26 TAL u 261 u 20 I-104 1 1 JAT 63) 1 -